Use the LCD to determine which fraction is greater
26. 2/5,3/10
Any help with this equation would be greatly appreciated:)

Answer by stanbon(75887)

Use the LCD to determine which fraction is greater
26. 2/5,3/10
LCD of the two fractions is 10
2/5 = 4/10
3/10 = 3/10
----------------
2/5 is greater than 3/10
=============
Cheers,
Stan H.